Please help me with this question.
X 1 3 6
Y 3 9 t
The table above shows a relation of the form y = kx` + c where k and c are constants.
calculate the value of t.

Answer by ankor@dixie-net.com(12701)

X 1 3 6
Y 3 9 t
The table above shows a relation of the form y = kx` + c where k and c are constants.
calculate the value of t.
:
Find k
The equation for x,y coordinates 1,3
3 = 1k + c
:
The equation for 3,9
9 = 3k + c
:
3k + c = 9
1k + c = 3
----------- subtract, eliminate c
2k = 6
k = 3
:
Find c:
3(3) + c = 9
c = 0
:
Solve for t, x = 6
t = 3x + 0
t = 3(6) + 0
t = 18